OpenStudy (firejay5):

Solve each equation. Then check your solution. Show and explain YOUR WORK so I can better understand it.

OpenStudy (firejay5):

\[\frac{ x }{ 2 } - \frac{ 1 }{ 3 } = \frac{ x }{ 3 } - \frac{ 1 }{ 2 }\]

OpenStudy (vuriffy):

Hmm. Well, I would say you must get x on one side.

OpenStudy (firejay5):

I know you have to get common denominators but I won't understand it unless it's written down because I am a visual learner

OpenStudy (vuriffy):

So, you may subtract x/2 then add 1/2 from the right side. Okay here:

OpenStudy (518nad):

-1 :)

OpenStudy (vuriffy):